OpenStudy (anonymous):

plz help asap Solve for x: 3/4 x + 5/8 = 4x

OpenStudy (campbell_st):

collect the line terms by subtracting 3/4x from both sides of the equation 5/8 = 3 3/4 x multiply both sides of the equation by 8 5 = 30x now solve for x

OpenStudy (anonymous):

so what would be the answer in the form

OpenStudy (campbell_st):

well divide both sides of the equation by 30 to fnd x. you would probably have to simplify the fraction

myininaya (myininaya):

you need x=a number he wants you to solve 5=30x but i think 5=30x isn't the right equation to be solving

OpenStudy (mathstudent55):

@campbell_st 3/4 x + 5/8 = 4x 5/8 = 3 1/4 x

OpenStudy (anonymous):

okay the answer is